# Unleashing Expert Opinion from Social Media for Stock Prediction

### v1

@@ III Problem Formulation
    Historical market data: For each stock $u\in\mathcal{S}$, we use price and fundamental features $X_{u,t}=[x_{u,t-L},\ldots,x_{u,t-1}]\in\mathbb{R}^{F\times L}$ from day $t-L$ to $t-1$ with a rolling window of size $L$. Unlike most stock prediction papers that use data up to timestep $t$, we use price features until $t-1$ to avoid look-ahead bias, as the exact price for day $t$ are only available after market close.
@@
    Expert signals: For each stock $u\in\mathcal{S}$, we perform real-time collection of StockTwits posts during day $t$. The collection process ends 5 minutes before market close, from which we extract expert signal $e_{u,t}$ if available.
@@
    Dynamic stock graph: To capture the relationships between stocks, we construct a dynamic graph $G_{t}=(V_{t},E_{t})$ based on price data up to day $t-1$, where nodes $V_{t}=\mathcal{S}$ represent the complete set of stocks and edges $E_{t}$ capture their relationships (see Sec. V.C Dual Graph Attention Network for details).
@@ IV Expert User Identification
Despite the presence of noise in social media data, there is still potential to harness it by developing more sophisticated models. One promising avenue is through building an expert tracing system, as proposed by several studies that attempted to identify expert users within social media platforms. Among them, the MFN model proposed by Wang et al.[20], the Smart-User-Classification model by Coyne et al. [19], the Winners model by Liao et al. [28] and “consistently correct user identification” by Bouadjenek et al.[27] aimed to classify expert users from social media posts that carry subjective sentiments. While the former three models rely on external sentiment analysis models to label tweets, which might introduce errors, our dataset from StockTwits contains user-labeled sentiments (bullish or bearish), which eliminates the need for external sentiment inference and reduces potential errors. However, when applied to our comprehensive dataset spanning the past five years, these methods all failed to identify experts/inverse experts who could consistently make (in)accurate predictions during the test period. Users identified as experts in the training phase did not perform reliably as experts in the test phase.
@@
    Sufficient Past Data Criteria: Existing methods employ insufficient criteria for expert identification, either using a fixed number of $N$ past posts (e.g., $N$=10) [28] or considering prediction accuracy only over a short period of $K$ days (e.g., $K$=40 or 90) [27, 20]. However, we found these methods unreliable. In our dataset, even after filtering out spammers, users who achieved over 80% accuracy in their past 20 posts had an average accuracy of only 54.0%, and users who achieved 80% accuracy in the past 40 days had an average accuracy of only 53.9% during the testing phase. Further analysis revealed that most users only have high prediction accuracy during specific periods, after which their accuracy declines significantly, and vice versa. For example, users who posted for more than 20 days in 2022 with an accuracy above 80% had an average accuracy about 50% in 2023, while users who had accuracy below 20% in 2022 saw their accuracy rise to nearly 50% in 2023. To better understand the shortcomings of existing approaches and address these limitations in expert identification, we propose two key criteria:
@@
    • Precise prediction across different market regimes: Some users consistently post only bullish predictions every day, while others post only bearish predictions. For those posting only bullish predictions, when the stock is in an upward trend, they may appear to be experts due to their constant bullish posts, but during market corrections, these same bullish posts lead to poor predictions, effectively turning them into inverse experts. Similarly, users who consistently post bearish predictions may seem expert-like during market downturns but become inverse experts during upward trends. These users have little value for analysis as they do not base their predictions on real market insight. To address this issue, not only do we consider the accuracy of the past $N$ tweets to ensure the dynamism of the experts, but we also need to look at their long-term predictions (two years) and see if they maintain a high level of accuracy, as most stocks will not remain in the same trend over an extended period of time.
@@
    • Focused Expert Evaluation: True market experts typically specialize in a manageable number of stocks to maintain deep analysis and insight. However, some users post multiple bullish or bearish sentiments for a large number of stocks on the same day. If the market happens to move in their predicted direction, they will appear highly accurate and be incorrectly identified as experts in previously mentioned expert identification methods. Such broad, unfocused posting behavior contradicts the nature of genuine expertise, which requires concentrated attention and thorough analysis. To ensure we identify truly focused experts, we require that a user’s past 20 posts must span at least five different trading days 111we tested various parameters and found that five different days yielded the best performance considering both the coverage and accuracy. This helps filter out users who spread their predictions too thin across many stocks without demonstrating sustained, focused analysis.
@@
By utilizing this approach, we developed a dynamic expert tracing system. For each trading day, if there is an expert whose past predictions meet the aforementioned criteria, we treat that expert’s prediction for the day as a representative trend signal, serving as the pseudo ground truth for whether the stock will rise or fall.
The procedure is summarized in Algorithm  1. For each trading day $d$, we first filter the posts to retain only the latest post per user-stock pair. For users who posted on day $d$, we evaluate their prediction accuracy over two time horizons. Let $T_{i,recent}$ and $F_{i,recent}$ represent the number of correct and incorrect predictions in their most recent $N$ posts, while $T_{i,long}$ and $F_{i,long}$ denote these numbers over the past $T$ years. To ensure both recent performance and consistent long-term accuracy, we require experts to meet two criteria: their most recent $N$ ($N$=20) posts must span at least $K$ ($K$=5) different days and achieve an accuracy above $P_{2}$ (80%), while maintaining a long-term accuracy above $P_{1}$ (65%) over the past $T$ ($T$=2) years. Similarly, inverse experts are identified when their recent and long-term accuracies fall below 1-$P_{2}$ (20%) and 1-$P_{1}$ (35%) respectively.
The expert tracing system achieved an accuracy rate as high as 77.26%. Furthermore, the inference stage has a time complexity of $O(len(C_{d}))$ where $\text{len}(C_{d})$ is the number of posts on the given day d. Since our Stocktwits database contains the historical posting data and accuracy for each user, the inference phase simply involves traversing all posts once to update the relevant information for each user. Therefore, the time taken to complete this operation scales linearly with the total number of posts. The low complexity, combined with the high accuracy, demonstrates the effectiveness and efficiency of our method in enhancing stock prediction models by filtering out noise and focusing on both expert and inverse expert insights.
@@
0:  current date $d$, minimum post requirement $N$, minimum days span $K$, tweet set $C$ with $C_{d}\subset C$ denoting tweets posted on day $d$, long-term accuracy threshold $P_{1}$, short-term accuracy threshold $P_{2}$, long-term range $T$
@@
1:  $C_{d}\leftarrow\text{filter\_daily\_latest}(C_{d})$Keep only the latest post per user-stock pair on day d}
2:  $I_{d}\leftarrow\text{all users of }C_{d}$
3:  for each $i\in I_{d}$ do
@@
4:     $C_{i,recent}\leftarrow\text{get\_recent\_N\_posts}(C,i,N)$Get most recent N posts for user $i$}
5:     $days_{i,recent}\leftarrow\text{count\_unique\_days}(C_{i,recent})$
6:     if $days_{i,recent}<K$ then
7:        continue
8:     end if
9:     $T_{i,recent}\leftarrow 0$; $F_{i,recent}\leftarrow 0$Recent N posts accuracy counters}
10:     for each $c\in C_{i,recent}$ do
11:        Check stock price change $\rho$ on the next trading day
12:        if $(c\in\text{bullish}\land\rho\in\text{rise})\lor(c\in\text{bearish}\land\rho\in\text{fall})$ then
13:           $T_{i,recent}\leftarrow T_{i,recent}+1$
14:        else
15:           $F_{i,recent}\leftarrow F_{i,recent}+1$
16:        end if
17:     end for
18:     $A_{i,recent}\leftarrow T_{i,recent}/(T_{i,recent}+F_{i,recent})$
19:     $C_{i,long}\leftarrow\text{get\_posts}(C_{d-T:d},i)$Get all posts in past T years}
20:     $T_{i,long}\leftarrow 0$; $F_{i,long}\leftarrow 0$Long-term accuracy counters}
21:     for each $c\in C_{i,long}$ do
22:        Check stock price change $\rho$ on the next trading day
23:        if $(c\in\text{bullish}\land\rho\in\text{rise})\lor(c\in\text{bearish}\land\rho\in\text{fall})$ then
24:           $T_{i,long}\leftarrow T_{i,long}+1$
25:        else
26:           $F_{i,long}\leftarrow F_{i,long}+1$
27:        end if
28:     end for
29:     $A_{i,long}\leftarrow T_{i,long}/(T_{i,long}+F_{i,long})$
30:     if $A_{i,recent}\geq P_{2}$ and $A_{i,long}\geq P_{1}$ then
@@
31:        Mark $i$ as expert and take long (short) positions for bullish (bearish) sentiment
32:     else if $A_{i,recent}\leq(1-P_{2})$ and $A_{i,long}\leq(1-P_{1})$ then
@@
33:        Mark $i$ as inverse expert and take short (long) positions for bullish (bearish) sentiment
34:     end if
35:  end for
@@ V Expert Opinion Propagation
Despite integrating both expert and inverse expert signals, like other expert identification approaches [20, 19, 27], our model still suffers from signal sparsity, producing meaningful predictions for only a few stocks on certain trading days. Although these predictions serve as a pseudo ground truth for stock trends, they account for only about 4% of our dataset, leaving most data points unlabeled. To address this limitation, we propose a graph-based method that transforms these limited expert predictions into practical signals and propagates them across related stocks, thereby extending coverage to inform all data points in the dataset.
@@ V-A Trend Signals Transformation
Though predicting whether a stock will rise or fall is important, the ultimate goal is not simply trend forecasting but maximizing excess returns through investment strategies. Feng et al. [2] has pointed out that traditional stock prediction methods focus either on classification (predicting the direction) or regression (predicting the price) are suboptimal because they fail to align with the goal of selecting stocks with the highest potential returns, which is the real objective for investors. Given that a more practical approach to stock prediction involves optimizing return ratios rather than focusing solely on trend direction, it becomes necessary to transform the pseudo ground truth from binary trend indicators (up or down) into continuous return ratios which can be treated as expert signal.
The transformation is as follows: for stocks predicted by experts to rise, we compute the average return ratio from days within the past 30 days when the stock showed an upward trend. Similarly, for a predicted downward trend, we use the average of the down days’ return ratios within the same period. This 30-day window for the return ratio which acts as a monthly indicator is a common approach in financial modelling for stock prediction [57, 58, 59] as it helps to smooth out daily volatility while providing a sufficient sample size to capture recent market trends. This calculated average thus serves as a ‘prior’ for return ratios, which effectively functions as the expert signal. If a stock is not predicted by experts on any given day, then the ‘prior’ for that stock is set to 0.
@@ V-B Temporal Pre-training Model
Now that we have obtained the sparse prior for the return ratios, the challenge remains that for over 95% of the stock-day pairs, we still lack any prior information for these pairs as they have no expert signal. To address this, we propose an initial step to train a temporal model, which will serve as a baseline for generating return ratio estimates. This step acts similarly to a pre-training process, where the temporal model captures patterns in stock movements without direct reliance on expert signals.
The model is trained using historical stock price data (open, high, low, close prices, and trading volume) and fundamental data. Unlike many studies that follow Feng et al.’s [2, 8, 53] method of normalizing the entire dataset across the whole time dimension (including test sets), which may lead to test set leakage, we normalize each feature daily across the stock cross-section.
@@ V-B1 MS-LSTM
Inspired by MS-RNN [60], we propose the pre-training model, termed MS-LSTM (Multi-Scale Long Short-Term Memory), which uses a multi-scale LSTM architecture. Each scale is handled by an independent LSTM network, allowing the model to capture temporal dependencies at varying resolutions. The trained model generates a foundational return ratio estimate for each stock daily, which acts as a baseline for stocks that lack expert signals.
Given an input sequence $\mathbf{X}\in\mathbb{R}^{N\times L\times d}$, where $N$ is the number of stocks, $L$ is the sequence length, and $d$ is the feature dimension, MS-LSTM processes the data through the following steps:

### v2

@@ III Problem Formulation
    Historical market data: For each stock $u\in\mathcal{S}$, we use price and fundamental features $X_{u,t}=[x_{u,t-L},\ldots,x_{u,t-1}]\in\mathbb{R}^{F\times L}$ from day tLt-L to t1t-1 with a rolling window of size $L$. Unlike most stock prediction papers that use data up to timestep $t$, we use price features until t1t-1 to avoid look-ahead bias, as the exact price for day $t$ are only available after market close.
@@
    Expert signals: For each stock $u\in\mathcal{S}$, we perform real-time collection of StockTwits posts during day $t$. The collection process ends 5 minutes before market close, from which we extract expert signals $e_{u,t}$ if available.
@@
    Dynamic stock graph: To capture the relationships between stocks, we construct a dynamic graph $G_{t}=(V_{t},E_{t})$ based on price data up to day t1t-1, where nodes $V_{t}=\mathcal{S}$ represent the complete set of stocks and edges $E_{t}$ capture their relationships (see Sec. V.C Dual Graph Attention Network for details).
@@ IV Expert User Identification
Despite the presence of noise in social media data, there is still potential to harness it by developing more sophisticated models. One promising avenue is through building an expert tracing system, as proposed by several studies that attempted to identify expert users within social media platforms. Among them, the MFN model proposed by Wang et al.[aaai_expert_based], the Smart-User-Classification model by Coyne et al. [Forecating_stock_prices], the Winners model by Liao et al. [liao2014winning] and “consistently correct user identification” by Bouadjenek et al.[A_user_centric] aimed to classify expert users from social media posts that carry subjective sentiments. While the former three models rely on external sentiment analysis models to label tweets, which might introduce errors, our dataset from StockTwits contains user-labeled sentiments (bullish or bearish), which eliminates the need for external sentiment inference and reduces potential errors. However, when applied to our comprehensive dataset spanning the past five years, these methods all failed to identify experts/inverse experts who could consistently make (in)accurate predictions during the test period. Users identified as experts in the training phase did not perform reliably as experts in the test phase.
@@
    Sufficient Past Data Criteria: Existing methods use insufficient criteria for expert identification, either using a fixed number of $N$ past posts (e.g., NN=10) [liao2014winning] or considering prediction accuracy only over a short period of $K$ days (e.g., KK=40 or 90) [A_user_centric, aaai_expert_based]. However, these methods proved unreliable. In our dataset, users who achieved over 80% accuracy in their past 20 posts had an average accuracy of only 54.0%, and users who achieved 80% accuracy in the past 40 days had an average accuracy of only 53.9% during the testing phase. Further analysis revealed that most users only have high prediction accuracy during specific periods, after which their accuracy declines significantly, and vice versa. For example, users who posted for more than 20 days in 2022 with an accuracy above 80% averaged about 50% in 2023, while users who had accuracy below 20% in 2022 saw their accuracy rise to nearly 50% in 2023. To better understand the shortcomings of existing approaches and address these limitations in expert identification, we propose two key criteria:
@@
    • Precise prediction across different market regimes: Some users consistently post only bullish predictions, while others post only bearish predictions. Those posting only bullish predictions may appear expert-like during upward trends but become inverse experts during market corrections. Similarly, consistently bearish users may seem expert-like during downturns but become inverse experts during upward trends. These users have little analytical value as they lack real market insight. To address this, we not only consider the accuracy of the past $N$ tweets to ensure expert dynamism, but also evaluate their long-term predictions over two years to verify sustained accuracy across market cycles. While existing expert identification methods often rely on short evaluation horizons (e.g., 40 or 90 days), this approach is vulnerable to market cycles. This 2-year evaluation window is grounded in the seminal work of Jegadeesh & Titman [jegadeesh1993returns], which demonstrates that strategies based on high 3- to 12-month past returns may experience significant reversals in the following year. Furthermore, [carhart1997persistence, fama2010luck] also demonstrate that longer evaluation horizons are essential, as users appearing expert-like during short-term momentum phases can become inverse experts when market regimes shift.
@@
    • Focused Expert Evaluation: True market experts typically specialize in a manageable number of stocks to maintain deep analysis and insight. However, some users post multiple bullish or bearish sentiments for numerous stocks on the same day. If the market happens to move in their predicted direction, they will appear highly accurate and be incorrectly identified as experts in previously mentioned expert identification methods. Such widespread, unfocused posting behavior contradicts the nature of genuine expertise, which requires concentrated attention and thorough analysis. This principle is grounded in cognitive science and financial research. Cognitive Load Theory recognizes the inherent limits on human information processing capacity [miller1956magical]. Deep financial analysis is cognitively demanding, and studies show that specialization correlates with superior forecast accuracy. For example, Clement [clement1999analyst] found that analysts with less complex portfolios (covering fewer firms) produce more accurate earnings forecasts. To ensure we identify truly focused experts, we require that a user’s past 20 posts must span at least five different trading days 111we tested various parameters and found that five different days yielded the best performance considering both the coverage and accuracy. This filters out users who employ a “shotgun” approach by spreading predictions across many stocks without demonstrating sustained analysis.
@@
By utilizing this approach, we developed a dynamic expert tracing system. For each trading day, if there is an expert whose past predictions meet the aforementioned criteria, we treat that expert’s prediction for the day as a representative trend signal, serving as the directional indicator for whether the stock will rise or fall.
@@
0: current date $d$, minimum post requirement NN, minimum days span requirement $K$, long-term accuracy threshold $P_{1}$, recent accuracy threshold $P_{2}$, long-term range $T$, tweet set $C$ with $C_{d}\subset C$ denotes tweets posted on day $d$, $C^{i}\subset C$ denotes all tweets by user $i$
@@
[…]
3:for each user $i\in I_{d}$ do
4:  {Stage 1: Evaluate Recent Performance}
@@
5:  $C^{i}_{\text{recent}}\leftarrow\text{get\_recent\_N\_posts}(C,i,N)$ $C^{i}_{\text{recent}}$ denotes the most recent N posts before day $d$ for user ii}
6:  $days^{i}_{recent}\leftarrow\text{count\_unique\_days}(C^{i}_{recent})$
7:  if $days^{i}_{recent}<K$ then
8:   continue
9:  end if
10:  $T^{i}_{recent}\leftarrow 0$; $F^{i}_{recent}\leftarrow 0$ {Recent accuracy counters}
11:  for each $c\in C^{i}_{recent}$ do
12:   Check stock price change $\rho$ on the next trading day
13:   if $(c\in\text{bullish}\land\rho\in\text{rise})\lor(c\in\text{bearish}\land\rho\in\text{fall})$ then
14:    $T^{i}_{recent}\leftarrow T^{i}_{recent}+1$
15:   else
16:    $F^{i}_{recent}\leftarrow F^{i}_{recent}+1$
17:   end if
18:  end for
19:  $A^{i}_{recent}\leftarrow T^{i}_{recent}/(T^{i}_{recent}+F^{i}_{recent})$
20:  {Stage 2: Evaluate Long-Term Performance}
21:  $C^{i}_{long}\leftarrow\text{get\_posts}(C_{d-T:d},i)$ {Get all posts in past T years}
22:  $T^{i}_{long}\leftarrow 0$; $F^{i}_{long}\leftarrow 0$ {Long-term accuracy counters}
23:  for each $c\in C^{i}_{long}$ do
24:   Check stock price change $\rho$ on the next trading day
25:   if $(c\in\text{bullish}\land\rho\in\text{rise})\lor(c\in\text{bearish}\land\rho\in\text{fall})$ then
26:    $T^{i}_{long}\leftarrow T^{i}_{long}+1$
27:   else
28:    $F^{i}_{long}\leftarrow F^{i}_{long}+1$
29:   end if
30:  end for
31:  $A^{i}_{long}\leftarrow T^{i}_{long}/(T^{i}_{long}+F_{i,long})$
32:  {Stage 3: Classify User}
33:  if $A^{i}_{recent}\geq P_{2}$ and $A^{i}_{long}\geq P_{1}$ then
@@
34:   Mark $i$ as expert and take long (short) positions for bullish (bearish) sentiment
35:  else if $A^{i}_{recent}\leq(1-P_{2})$ and $A^{i}_{long}\leq(1-P_{1})$ then
@@
36:   Mark $i$ as inverse expert and take short (long) positions for bullish (bearish) sentiment
37:  end if
38:end for
Our dynamic expert tracing system, summarized in Algorithm 1, identifies experts for any given trading day $d$. The system relies on a historical database of all user posts. The algorithm first filters to keep only the latest post per user-stock pair on day $d$. It then iterates through each user who posted that day. For each user $i$, the system performs a two-stage evaluation using their historical posting data.
First, in the recent performance evaluation stage, we assess the user’s most recent $N$ posts ($N=20$). To ensure the user is active and focused, we require these posts to span at least $K$ days ($K=5$). We calculate $T^{i}_{recent}$ and $F^{i}_{recent}$, representing the number of correct and incorrect predictions in these recent posts. This gives us their recent accuracy, $A^{i}_{recent}$. Second, the long-term performance evaluation stage assesses the user’s track record over the past $T=2$ years ($T=2$). We compute $T^{i}_{long}$ and $F^{i}_{long}$, the total counts of correct and incorrect predictions over this period, to determine their long-term accuracy $A^{i}_{long}$. Finally, a user is identified as an expert if they meet two criteria: their recent accuracy $A^{i}_{recent}$ must exceed threshold $P_{2}$ (80%), and their long-term accuracy $A^{i}_{long}$ must exceed threshold $P_{1}$ (65%). Similarly, a user is classified as an inverse expert if $A^{i}_{recent}$ falls below $1-P_{2}$ (20%) and $A^{i}_{long}$ is below $1-P_{1}$ (35%). The rationale for these threshold choices and sensitivity analysis can be found in Supplementary Material S1.
The expert tracing system achieved an accuracy rate as high as 72.8%. Furthermore, the inference stage has time complexity $O(\text{len}(C_{d}))$, where $\text{len}(C_{d})$ is the number of posts on day $d$. Since our StockTwits database contains historical posting data and accuracy for each user, the inference phase simply traverses all posts once to update relevant user information. The operation scales linearly with total posts. The low complexity and high accuracy demonstrate our method’s effectiveness in enhancing stock prediction by filtering noise and focusing on expert and inverse expert insights.
@@ V Expert Opinion Propagation
Despite integrating both expert and inverse expert signals, like other expert identification approaches [aaai_expert_based, Forecating_stock_prices, A_user_centric], our model still suffers from signal sparsity. These signals produce meaningful predictions for only few stocks on some trading days, covering merely 4% of our dataset. Consequently, a strategy built directly on these sparse signals would suffer from high exposure and excessive concentration risk. To address this limitation, we propose a graph-based method that transforms these limited expert predictions into practical signals and propagates them across related stocks, thereby extending coverage to inform all data points in the dataset.
@@ V-A Trend Signals Transformation
Though predicting whether a stock will rise or fall is important, the ultimate goal is not simply trend forecasting but maximizing excess returns through investment strategies. Feng et al. [feng2019temporal] pointed out that traditional stock prediction methods focusing on classification (predicting direction) or regression (predicting price) are suboptimal because they fail to align with selecting stocks with the highest potential returns, which is the real objective for investors. Given that a more practical approach involves optimizing return ratios rather than focusing solely on trend direction, it becomes necessary to transform the expert prediction which is the binary trend indicators into continuous return ratios which can be treated as expert signals.
The transformation is as follows: for stocks predicted by experts to rise, we compute the average return ratio from days within the past 30 days when the stock showed an upward trend. Similarly, for a predicted downward trend, we use the average of the down days’ return ratios within the same period. This 30-day window for the return ratio which acts as a monthly indicator is a common approach in financial modelling for stock prediction [Forecasting_significant_stock, Stock_market_prediction, A_systematic_review] as it helps to smooth out daily volatility while providing a sufficient sample size to capture recent market trends. We use this calculated average as the expert signal. If a stock is not predicted by experts on any given day, then the expert signal for that stock is set to 0. We compare this trend signals transformation approach against simply using binary signals (+1/-1), with experimental results demonstrating the superiority of our method (see Supplementary Material S2).
@@ V-B Temporal Pre-training Model
Now that we have obtained the sparse expert signals, the challenge remains that for over 96% of stock-day pairs, we still lack any expert prediction for these pairs. To address this, we propose an initial step to train a temporal model, which will serve as a baseline for generating return ratio estimates. This step acts similarly to a pre-training process, where the temporal model captures patterns in stock movements without direct reliance on expert signals.
The model is trained using historical stock price data (open, high, low, close prices, and trading volume) and fundamental data. Unlike many studies following Feng et al.’s [feng2019temporal, Deep_attentive_learning, wang2022hatr] method of normalizing the entire dataset across the whole time dimension (including test sets), which may lead to test set leakage, we normalize each feature daily across the stock cross-section.
@@ V-B1 MS-LSTM
Inspired by MS-RNN [chung2022hierarchical], we propose the pre-training model, termed MS-LSTM (Multi-Scale Long Short-Term Memory). Each scale is handled by an independent LSTM network, allowing the model to capture temporal dependencies at varying resolutions. The trained model generates a foundational return ratio estimate for each stock daily, which acts as a baseline for stocks that lack expert signals.
Given input sequence $\mathbf{X}\in\mathbb{R}^{N\times L\times d}$, where $N$ is the number of stocks, $L$ is sequence length, and $d$ is feature dimension, MS-LSTM processes the data as follows: